Compare the relativistic momentum of an electron to its classical momentum when it has a speed of (a) 0.88c in an accelerator used for treating cancer and (b)4.50 × 107 𝑚/𝑠 in the cathode ray tube of a TV set.






1
Expert's answer
2022-06-02T10:25:39-0400
"p_c=mv,\; p_r=\frac{mv}{\sqrt{1-(v/c)^2}}"

(a)

"p_c=9.1*10^{-31}*0.88*3*10^8\\=2.4*10^{-22}\:{\rm kg\cdot m/s}"

"p_r=\frac{mv}{\sqrt{1-(v/c)^2}}"

"=\frac{2.4*10^{-22}\:{\rm kg\cdot m/s} }{\sqrt{1-(0.88)^2}}=5.1*10^{-22}\:{\rm kg\cdot m/s}"

(b)

"p_c=9.1*10^{-31}*4.5*10^7\\=4.1*10^{-23}\:{\rm kg\cdot m/s}"

"p_r=\frac{mv}{\sqrt{1-(v/c)^2}}"

"=\frac{4.1*10^{-23}\:{\rm kg\cdot m/s} }{\sqrt{1-(4.5*10^7/3*10^8)^2}}=4.1*10^{-23}\:{\rm kg\cdot m/s}"